How many heartbeats in a typical human lifetime?
1
Expert's answer
2020-06-23T13:07:03-0400

As per the question, we have to find the number of heartbeat of a typical human in his lifetime,

As be know that 72 heart beat in 1 min,

so in one day total 24 hours, So number of heartbeat in one day "=24\times 60\times 74"

number of heartbeat in one year "=24\times60 \times 75\times365"

total number of years in one typical human lifetime="75 years"

Hence total number of heartbeat in the lifetime of the typical human "=24\times 60\times 75\times 365=39420000"

"=3.942\times 10^{7}"